From the give table calculate Elasticity of Price, Total Revenue and Marginal Revenue. Also, explain the relationship between AR and MR?

Perice

6

5

4

Quantity


100

200

TR & MR




1
Expert's answer
2021-04-12T07:08:13-0400

P Q TR MR Ed

6 0 0 - -

5 100 500 500 -11

4 200 800 300 -3

AR is average revenue (TR/Q) and MR is marginal revenue (MR = TR'(Q)). TR = P×Q.

"Ed = \frac{Q2 - Q1} {P2 - P1} ×\frac{P2 + P1} {Q2 + Q1}."